OpenStudy (anonymous):

na please help me with this one to find the solution of the differential equation [(x+1)^2]d^2y/dx^2 + (x+1)dy/dx + y = cos[ln(x+1)]

OpenStudy (loser66):

Homogeneous part: Let t = x+1 --> \(t^2y''+ty'+y =0\) It's an Euler form. Characteristic equation: r^2+1 =0 gives us complex root \(r=\pm i\) so, general solution is \(y_c = C_1cos (lnt)+C_2sin (lnt)\) replace t = x+1, we have \(y_c = C_1cos(ln (x+1) +C_2sin(ln(x+1)\) Partial part: y = t cos (lnt) I think you can step up, right?

OpenStudy (anonymous):

yep i can but where does the euler equation form from

OpenStudy (loser66):

I am sorry, it's indicial equation, not characteristic equation

OpenStudy (anonymous):

and how is r^2 + 1= 0 the complimentary part's equation giving function

OpenStudy (loser66):

From your original one , after let t = x+1, we get that form, just apply formula

OpenStudy (loser66):

r^2+1 =0 ---> r^2 =-1 ---> r = \(\pm i\)

OpenStudy (anonymous):

yea i get that but why will r^2 give us the general solution, where does that come from

OpenStudy (loser66):

That's the step to solve Euler's ODE

OpenStudy (anonymous):

yep i get the t^2y'' + ty+ +y = 0 part is by substution, where does the r^2+1 come from

OpenStudy (anonymous):

ahh (Y) i'll check it out

OpenStudy (anonymous):

how/why is cos{ln(x+1)} equated to zero

OpenStudy (anonymous):

ahahahahaha sorry na to bug you that much ^3^)7 i'll just check it out myself

OpenStudy (loser66):

I don't let it = 0. just separate to solve homogeneous part and partial part.

OpenStudy (anonymous):

so the substution gives the homogenous part?

OpenStudy (loser66):

for convenient, you should substitute to both.

OpenStudy (anonymous):

i shall (Y)

OpenStudy (loser66):

what does (Y) mean in "i shall (Y)"?

OpenStudy (anonymous):

it means the thumbs up sign on fb XD so a bit of habit,a bad one i guess